How much money would you have to deposit today in order to have $2,000 in three years if the discount rate is 6 percent per year? (Do not round intermediate calculations. Round your answer to 2 decimal places. (e.g., 32.16))

what is the amonunt of deposit ??
Answer #1

Amount to be deposited today is the present value of $2,000 receivable three years from now discounted at 6%

Present value factor

= 1 / (1 + r) ^ n

Where,

r = Rate of interest = 6% or 0.06

n = Years = 3

So, Present value factor

= 1 / (1.06^3)

= 1 / 1.191016

= 0.839619

So, present value

= Future value x Present value factor

= $2,000 x 0.839619

= $ 1,679.24